### v1

# An averaging formula for Nielsen numbers of affine $n$valued maps on infra-nilmanifolds

Abstract.
In [8, 9], the authors developed a nice formula to compute the Nielsen number of a self-map on an infra-nilmanifold. For the case of nilmanifolds this formula was extended to $n$-valued maps in [4]. In this paper, we extend these results further and establish the averaging formula to compute the Nielsen number of any $n$-valued affine map on an infra-nilmanifold.
Research supported by Methusalem grant METH/21/03 – long term structural funding of the Flemish Government.
Introduction
Let $M$ be a closed manifold (or more generally a compact polyhedron) and let $f:M\to M$ be a self-map of $M$. In Nielsen-Reidemeister fixed point theory one attaches to each such map $f$ a number $N(f)$, the Nielsen number of $f$, which is a (sharp) lower bound for the number of fixed points of any map $g$ which is homotopic to $f$. In this paper we are concerned with infra-nilmanifolds. It is known that any self-map on an infra-nilmanifold is homotopic to a so-called affine map ([10]). These maps are rather algebraic in nature and one has developed an averaging formula for computing their Nielsen number ([8, 9]). Since the Nielsen number of a map is invariant under a homotopy, this formula allows us to compute the Nielsen number of any self-map on an infra-nilmanifold. For $n$-valued maps the situation is more complicated. It is no longer true that any $n$-valued map on an infra-nilmanifold is homotopic to an affine $n$-valued map ([6]), but nevertheless, these affine $n$-valued maps still form a large and interesting class of $n$-valued maps on infra-nilmanifolds. In this paper, we obtain a natural generalisation of the formula from [9] for computing the Nielsen number of any affine $n$-valued map on an infra-nilmanifold.
In the first section we recall the basic Nielsen theory for $n$-valued maps and describe a general averaging formula (1.1) for computing the Nielsen number of an $n$-valued map. In the second section, we recall the necessary concepts concerning infra-nilmanifolds and affine $n$-valued maps. In the third section we establish a decomposition of the fixed point set of an $n$-valued map on an infra-nilmanifold, which allows us to rewrite the general averaging formula into a more algebraic one (3.1). In the fourth section we further examine this formula in the case of affine $n$-valued maps, to finally arrive at the main result of this paper (Theorem 4.9), which is an averaging formula generalising in a natural way the formula from [9].
1. Nielsen theory of $n$-valued maps
Let $X$ be a connected compact manifold with fundamental group $\pi$ and universal cover $p:\tilde{X}\to X$. We will view $\pi$ as being the group of covering transformations of $p$.
An n-valued map $f:X\multimap X$ is a continuous (i.e. upper and lower semi-continuous) set-valued function, sending each point of $X$ to a set of $n$ distinct points in $X$. Equivalently, by the work of Brown and Gonçalves [2], the map $f$ can be viewed as a continuous single-valued function from $X$ to the unordered configuration space
$D_{n}(X)=\{\{x_{1},\ldots,x_{n}\}\subseteq X\mid x_{i}\neq x_{j}\text{ if }i\neq j\},$
i.e. the quotient of the space $F_{n}(X)=\{(x_{1},\ldots,x_{n})\in X^{n}\mid x_{i}\neq x_{j}\text{ if }i\neq j\}$ under the action of the permutation group $\Sigma_{n}$. The fixed point set of an $n$-valued map $f$ is $\mathrm{Fix}(f)=\{x\in X\mid x\in f(x)\}$.
In what follows we recall some of the results of Brown et al. [1], where the study of fixed points of $f$ was done by considering fixed points of lifts of $f$ to the orbit configuration space
$F_{n}(\tilde{X},\pi)=\{(\tilde{x}_{1},\ldots,\tilde{x}_{n})\in\tilde{X}^{n}\mid p(\tilde{x}_{i})\neq p(\tilde{x}_{j})\text{ if }i\neq j\},$
which is a covering space of $D_{n}(X)$ with covering map
$p^{n}:F_{n}(\tilde{X},\pi)\to D_{n}(X):(\tilde{x}_{1},\ldots,\tilde{x}_{n})\mapsto\{p(\tilde{x}_{1}),\ldots,p(\tilde{x}_{n})\}.$
The corresponding covering group is $\pi^{n}\rtimes\Sigma_{n}$, which acts on $F_{n}(\tilde{X},\pi)$ by
$(\gamma_{1},\ldots,\gamma_{n};\sigma)(\tilde{x}_{1},\ldots,\tilde{x}_{n})=(\gamma_{1}\tilde{x}_{\sigma^{-1}(1)},\ldots,\gamma_{n}\tilde{x}_{\sigma^{-1}(n)}).$
A lift of $f$ can then be defined as a map $\tilde{f}:\tilde{X}\to F_{n}(\tilde{X},\pi)$ such that
[…]
commutes. As $F_{n}(\tilde{X},\pi)$ is a subspace of $\tilde{X}^{n}$, such a lift $\tilde{f}:\tilde{X}\to F_{n}(\tilde{X},\pi)$ splits into $n$ maps $\tilde{f}_{1},\ldots,\tilde{f}_{n}:\tilde{X}\to\tilde{X}$, called the lift-factors of $\tilde{f}$.
A lift $\tilde{f}=(\tilde{f}_{1},\ldots,\tilde{f}_{n})$ induces a morphism $(\phi_{1},\ldots,\phi_{n};\sigma):\pi\to\pi^{n}\rtimes\Sigma_{n}$ of the covering groups so that for all $\gamma\in\pi$,
$(\tilde{f}_{1}\gamma,\ldots,\tilde{f}_{n}\gamma)=(\phi_{1}(\gamma)\tilde{f}_{\sigma_{\gamma}^{-1}(1)},\ldots,\phi_{n}(\gamma)\tilde{f}_{\sigma_{\gamma}^{-1}(n)}).$
Using this morphism, one can first subdivide $\{1,\ldots,n\}$ into $\sigma$-classes, the equivalence classes for the relation given by
$i\sim j\enskip\Leftrightarrow\enskip\exists\gamma\in\pi:\sigma_{\gamma}(i)=j.$
For all $i$, denote by $S_{i}=\{\gamma\in\pi\mid\sigma_{\gamma}(i)=i\}$ the stabilizer of $i$ for this relation, which is a finite index subgroup of $\pi$. The restrictions of the maps $\phi_{i}$ to $S_{i}$ are group morphisms, using which one can define an equivalence relation on $\pi$ by
$\alpha\sim\beta\enskip\Leftrightarrow\enskip\exists\gamma\in S_{i}:\alpha=\gamma\beta\phi_{i}(\gamma)^{-1}.$
The set of equivalence classes for this relation is denoted $\mathcal{R}[\phi_{i}]$, and its number of elements is the Reidemeister number $R(\phi_{i})$.
Remark 1.1.
In general, given two group morphisms $\varphi,\psi:\Pi_{1}\to\Pi_{2}$, one can consider the equivalence relation on $\Pi_{2}$ given by
$\alpha\sim\beta\enskip\Leftrightarrow\enskip\exists\gamma\in\Pi_{1}:\alpha=\psi(\gamma)\beta\varphi(\gamma)^{-1}.$
The equivalence classes are called the Reidemeister coincidence classes of $\varphi$ and $\psi$, and denoted $\mathcal{R}[\varphi,\psi]$. In case $\Pi_{1}$ is a subgroup of $\Pi_{2}$ and $\psi$ is the inclusion $\iota:\Pi_{1}\to\Pi_{2}$, we abbreviate $\mathcal{R}[\varphi,\iota]=\mathcal{R}[\varphi]$.
The $\sigma$-classes and Reidemeister classes of $\phi_{i}$ partition the fixed point set of $f$ into a disjoint union of fixed point classes: if $i_{1},\ldots,i_{r}$ are representatives of the $\sigma$-classes, then
$\mathrm{Fix}(f)=\bigsqcup_{\ell=1}^{r}\bigsqcup_{[\alpha]\in\mathcal{R}[\phi_{i_{\ell}}]}p(\mathrm{Fix}(\alpha\tilde{f}_{i_{\ell}})).$
To each fixed point class one can associate an index, and the fixed point classes with non-zero index are called essential. The number of essential fixed point classes is the Nielsen number of $f$, which is a lower bound for the number of fixed points among all maps homotopic to $f$:
$N(f)\leq\text{min}\{\#\mathrm{Fix}(g)\mid g\sim f\}.$
By the above, the Nielsen number can be written as
$N(f)=\sum_{\ell=1}^{r}\sum_{[\alpha]\in\mathcal{R}[\phi_{i_{\ell}}]}\varepsilon_{\alpha,i_{\ell}}$
where
$\varepsilon_{\alpha,i}=\left\{\begin{array}[]{ll}1&\text{if $p(\mathrm{Fix}(\alpha\tilde{f}_{i}))$ is an essential fixed point class of $f$,}\\ 0&\text{otherwise.}\end{array}\right.$
Since the representatives for the $\sigma$-classes were chosen arbitrarily, and choosing another representative $i_{\ell}$ for some $\ell$ must give the same result for $N(f)$, the number
$\sum_{[\alpha]\in\mathcal{R}[\phi_{i_{\ell}}]}\varepsilon_{\alpha,i_{\ell}}$
does not depend on the chosen representative $i_{\ell}$. Therefore the formula above can also be written as
(1.1) $N(f)=\sum_{i=1}^{n}\frac{1}{[\pi:S_{i}]}\sum_{[\alpha]\in\mathcal{R}[\phi_{i}]}\varepsilon_{\alpha,i}.$
2. Infra-nilmanifolds and affine maps
The maps we consider in this paper are affine $n$-valued maps on infra-nilmanifolds. An infra-nilmanifold, as defined e.g. in [5], is the quotient of a connected and simply connected nilpotent Lie group $G$ under the action of an almost-Bieberbach group $\pi\subseteq G\rtimes\mathrm{Aut}(G)$, i.e. a torsion free cocompact discrete subgroup of $G\rtimes C$, with $C$ a maximal compact subgroup of $\mathrm{Aut}(G)$. Such an almost-Bieberbach group acts properly discontinuously on $G$ and hence $\pi$ is (isomorphic to) the fundamental group of the infra-nilmanifold $\pi\backslash G$. As before, we will view $\pi$ as the covering group of the universal covering $G\to\pi\backslash G$.
In the special case where this group $\pi$ is entirely contained in $G$, the quotient space is called a nilmanifold. Every infra-nilmanifold $\pi\backslash G$ is finitely covered by a nilmanifold $N\backslash G$: it follows from the generalised Bieberbach Theorems that $\pi\cap G$ is of finite index in $\pi$, so one can take $N=\pi\cap G$ or any other subgroup $N\subseteq\pi\cap G$ which is of finite index in $\pi$. Such a group $N$ is a finitely generated torsion free nilpotent group and is a lattice of the Lie group $G$.
Nilmanifolds have been studied in detail by Mal’cev; see [11]. We briefly recall some of his results.
For a nilpotent group $G$ of class $c$, we denote the lower central series by
$G=\gamma_{1}(G)\geq\ldots\geq\gamma_{c}(G)\geq\gamma_{c+1}(G)=1.$
For a finitely generated torsion free nilpotent group $N$ of class $c$ we consider another central series whose terms are the isolators of the lower central series,
$N_{i}=\sqrt[N]{\gamma_{i}(N)}=\{g\in N\mid\exists k>0:g^{k}\in\gamma_{i}(N)\}.$
This series has the useful property that for all $i$, the quotient $N_{i}/N_{i+1}$ is free abelian of finite rank, say $N_{i}/N_{i+1}\cong\mathbb{Z}^{k_{i}}$. Moreover, for these numbers $k_{i}$, the factors of the lower central series of $G$ satisfy $\gamma_{i}(G)/\gamma_{i+1}(G)\cong\mathbb{R}^{k_{i}}$. Morphisms of (lattices in) $G$ induce morphisms on these quotients, which can be represented by $k_{i}\times k_{i}$ real (or integer) matrices.
If $N,N^{\prime}$ are lattices in $G$, any morphism $\phi:N\to N^{\prime}$ extends uniquely to a Lie group morphism $G\to G$. The latter induces a morphism on the Lie algebra $\mathfrak{g}$ of $G$. From now on we will refer to this as the Lie algebra morphism induced by $\phi:N\to N^{\prime}$, and write $\phi_{*}:\mathfrak{g}\to\mathfrak{g}$, without explicitly mentioning the extension of $\phi$ to $G$.
Another useful property of connected and simply connected nilpotent Lie groups $G$ is that the exponential map $\exp:\mathfrak{g}\to G$ is a diffeomorphism [3, Theorem 1.2.1]. In particular, it is invertible, and we denote the inverse map by $\log:G\to\mathfrak{g}$. This allows us to define the continuous function
$G\times\mathbb{R}\to G:(g,t)\mapsto g^{t}=\exp(t\log(g))$
which we will use later on.
On an infra-nilmanifold we can define the notion of an affine $n$-valued map. A map $f:\pi\backslash G\to D_{n}(\pi\backslash G)$ is affine if it lifts to a map of the form
$\tilde{f}:G\to F_{n}(G,\pi):x\mapsto(g_{1}\varphi_{1}(x),\ldots,g_{n}\varphi_{n}(x))$
with $g_{1},\ldots,g_{n}\in G$ and $\varphi_{1},\ldots,\varphi_{n}\in\mathrm{End}(G)$. We remark that this condition is independent of the chosen lift $\tilde{f}$. The corresponding maps $\phi_{i}$ and $\sigma$ from section 1 satisfy
(2.1) $g_{i}\varphi_{i}(\gamma x)=\phi_{i}(\gamma)g_{\sigma_{\gamma}^{-1}(i)}\varphi_{\sigma_{\gamma}^{-1}(i)}(x)$
for all $\gamma\in\pi$, $x\in G$ and $i\in\{1,\ldots,n\}$.
Nielsen numbers of affine $n$-valued maps on nilmanifolds have been studied in [4], where the following formula is proven:
Theorem 2.1 ([4, Theorem 6.13]).
Let $N\backslash G$ be a nilmanifold. The Nielsen number of an affine map $f:N\backslash G\to D_{n}(N\backslash G)$ with lift $\tilde{f}:G\to F_{n}(G,N):x\mapsto(g_{1}\varphi_{1}(x),\ldots,g_{n}\varphi_{n}(x))$ is given by
$N(f)=\sum_{i=1}^{n}|\det(I-(\varphi_{i})_{*})|.$
The goal of this paper is to generalise this formula to infra-nilmanifolds. In the single-valued case, this extension from nilmanifolds to infra-nilmanifolds was done by Kim, Lee and Lee in [8]. They proved that any map $f$ on an infra-nilmanifold $\pi\backslash G$ can be lifted to a nilmanifold $N\backslash G$ that finitely covers $\pi\backslash G$, and they expressed the Nielsen number of $f$ as the average of the Nielsen numbers of these lifts.
We cannot follow the same approach in the $n$-valued case, since $n$-valued maps in general no longer admit lifts to a nilmanifold. However, we can decompose the formula for the Nielsen number in a similar way as in [8], and then plug in the knowledge about affine maps on nilmanifolds established in [4] to obtain the desired averaging formula without involving actual lifts to the nilmanifold.
3. Decomposition of the fixed point set
Let $\pi\backslash G$ be an infra-nilmanifold, finitely covered by a nilmanifold $N\backslash G$. For a general map $f:\pi\backslash G\to D_{n}(\pi\backslash G)$, we will decompose the sum (1.1) into a sum over Reidemeister classes in $N$ and in $\pi/N$, generalising the approach in [8].
For all $i$, consider the group $S^{\prime}_{i}$ generated by $\{\gamma^{[\pi:N]}\mid\gamma\in S_{i}\}$. This is a normal subgroup of $S_{i}$, and the quotient group $S_{i}/S^{\prime}_{i}$ is periodic, since the order of every element divides $[\pi:N]$. Also note that $\pi$ (and therefore also $S_{i}/S^{\prime}_{i}$) is polycyclic-by-finite, since its finite index normal subgroup $N$ is finitely generated and nilpotent. Since periodic polycyclic-by-finite groups are automatically finite, it follows that $[S_{i}:S^{\prime}_{i}]$ is finite. Consequently, also
$[N:S^{\prime}_{i}]=\frac{[\pi:S_{i}][S_{i}:S^{\prime}_{i}]}{[\pi:N]}$
is finite. Thus, $S_{i}^{\prime}$ is a finite index normal subgroup of $S_{i}$ and a finite index subgroup of $N$.
Furthermore, note that the image of $S^{\prime}_{i}$ under any morphism $S_{i}\to\pi$ is contained in $N$. Hence the inclusion map $\iota:S_{i}\to\pi$ and the morphism $\phi_{i}:S_{i}\to\pi$ induce morphisms
$\iota^{\prime}:S_{i}^{\prime}\to N$ $:\gamma\mapsto\gamma$ $\bar{\iota}:S_{i}/S_{i}^{\prime}\to\pi/N$ $:\gamma S_{i}^{\prime}\mapsto\gamma N$
$\phi^{\prime}_{i}:S_{i}^{\prime}\to N$ $:\gamma\mapsto\phi_{i}(\gamma)$ $\bar{\phi}_{i}:S_{i}/S_{i}^{\prime}\to\pi/N$ $:\gamma S_{i}^{\prime}\mapsto\phi_{i}(\gamma)N.$
We will decompose the set of Reidemeister classes of $\phi_{i}$ using Reidemeister classes associated to $\bar{\phi}_{i}$ and $\phi_{i}^{\prime}$.
First, note that
$\mathcal{R}[\phi_{i}]=\bigcup_{\bar{\alpha}\in\pi/N}\,\bigcup_{g\in N}\,[g\alpha].$
If $[g\alpha]=[g^{\prime}\alpha^{\prime}]$ in $\mathcal{R}[\phi_{i}]$, then certainly $[\bar{\alpha}]=[\bar{\alpha}^{\prime}]$ in $\mathcal{R}[\bar{\phi}_{i},\bar{\iota}]$, i.e.
$\exists\bar{\gamma}\in S_{i}/S^{\prime}_{i}:\bar{\alpha}=\bar{\iota}(\bar{\gamma})\bar{\alpha}^{\prime}\bar{\phi}_{i}(\bar{\gamma})^{-1}.$
Conversely, note that if $[\bar{\alpha}]=[\bar{\alpha}^{\prime}]$ in $\mathcal{R}[\bar{\phi}_{i},\bar{\iota}]$, then there is a $g\in N$ so that $[g\alpha]=[\alpha^{\prime}]$ in $\mathcal{R}[\phi_{i}]$.
For fixed $\alpha$, we have $[g\alpha]=[g^{\prime}\alpha]$ in $\mathcal{R}[\phi_{i}]$ if and only if $[g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$, with $\tau_{\alpha}:\pi\to\pi:\beta\mapsto\alpha\beta\alpha^{-1}$; i.e.
$\exists\gamma\in S_{i}:g=\gamma g^{\prime}\alpha\phi_{i}(\gamma)^{-1}\alpha^{-1}.$
Thus, the above are disjoint unions when taken over these respective Reidemeister classes:
$\mathcal{R}[\phi_{i}]=\bigsqcup_{[\bar{\alpha}]\in\mathcal{R}[\bar{\phi}_{i},\bar{\iota}]}\bigsqcup_{[g]\in\mathcal{R}[\tau_{\alpha}\phi_{i}]}\,[g\alpha].$
Rather than in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$, we would like to view the classes $[g]$ with $g\in N$ as elements of $\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$, where
$g\sim g^{\prime}\enskip\Leftrightarrow\enskip\exists\gamma\in S^{\prime}_{i}:g=\gamma g^{\prime}\alpha\phi^{\prime}_{i}(\gamma)^{-1}\alpha^{-1}.$
Since $S_{i}^{\prime}\subseteq S_{i}$, it is clear that $[g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ implies $[g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$, but conversely there can be distinct elements in $\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ that are equivalent in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$. The following lemma allows us to describe how many different classes in $\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ give rise to the same class in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$. In this lemma, we use $\mathrm{coin}(\varphi,\psi)=\{\gamma\in\Pi_{1}\mid\varphi(\gamma)=\psi(\gamma)\}$ to denote the group of coincidence points of two morphisms $\varphi,\psi:\Pi_{1}\to\Pi_{2}$.
Lemma 3.1 ([7, Lemma 3.1.4]).
Given a commutative diagram of groups with exact rows
[…]
such that the quotient groups $\Pi_{i}/\Gamma_{i}$ are finite. For all $\alpha\in\Pi_{2}$ and $g\in\Gamma_{2}$, the number of Reidemeister coincidence classes $[g^{\prime}]\in\mathcal{R}[\tau_{\alpha}\varphi^{\prime},\psi^{\prime}]$ such that $[g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\varphi,\psi]$ is equal to
$[\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\varphi},\bar{\psi}):\overline{\mathrm{coin}(\tau_{g\alpha}\varphi,\psi)}]$
where $\overline{\mathrm{coin}(\tau_{g\alpha}\varphi,\psi)}$ denotes the image of $\mathrm{coin}(\tau_{g\alpha}\varphi,\psi)$ in $\Pi_{1}/\Gamma_{1}$.
Applying this to the diagram
[…]
learns that for a given $g\in N$, the number of classes $[g^{\prime}]\in\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ such that $[g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$ is equal to
$[\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\phi}_{i},\bar{\iota}):\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}]=\frac{|\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\phi}_{i},\bar{\iota})|}{|\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}|},$
where $\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}$ denotes the image of $\mathrm{fix}(\tau_{g\alpha}\phi_{i})=\mathrm{coin}(\tau_{g\alpha}\phi_{i},\iota)$ in $S_{i}/S_{i}^{\prime}$. Therefore the sum (1.1) decomposes as
$N(f)=\sum_{i=1}^{n}\frac{1}{[\pi:S_{i}]}\sum_{[\bar{\alpha}]\in\mathcal{R}[\bar{\phi}_{i},\bar{\iota}]}\sum_{[g]\in\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]}\frac{|\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}|}{|\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\phi}_{i},\bar{\iota})|}\,\varepsilon_{g\alpha,i}.$
To further simplify this formula, consider the action of $S_{i}/S^{\prime}_{i}$ on $\pi/N$ given by
$\bar{\gamma}\cdot\bar{\alpha}=\bar{\iota}(\bar{\gamma})\bar{\alpha}\bar{\phi}_{i}(\bar{\gamma})^{-1}.$
The orbit of an element $\bar{\alpha}$ for this action is its Reidemeister class $[\bar{\alpha}]\in\mathcal{R}[\bar{\phi}_{i},\bar{\iota}]$. The stabilizer of $\bar{\alpha}$ is the set
$\{\bar{\gamma}\in S_{i}/S^{\prime}_{i}\mid\bar{\iota}(\bar{\gamma})\bar{\alpha}\bar{\phi}_{i}(\bar{\gamma})^{-1}=\bar{\alpha}\}=\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\phi}_{i},\bar{\iota}).$
Therefore we have
$|[\bar{\alpha}]|\cdot|\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\phi}_{i},\bar{\iota})|=|S_{i}/S^{\prime}_{i}|=[S_{i}:S^{\prime}_{i}].$
This reduces the expression for $N(f)$ to
$N(f)$ $=\sum_{i=1}^{n}\frac{1}{[\pi:S_{i}]}\sum_{\bar{\alpha}\in\pi/N}\frac{1}{|[\bar{\alpha}]|}\sum_{[g]\in\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]}\frac{|\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}|}{|\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\phi}_{i},\bar{\iota})|}\,\varepsilon_{g\alpha,i}$
(3.1) $=\sum_{i=1}^{n}\frac{1}{[\pi:S^{\prime}_{i}]}\sum_{\bar{\alpha}\in\pi/N}\sum_{[g]\in\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]}|\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}|\,\varepsilon_{g\alpha,i}.$
4. Averaging formula for affine maps
Now we apply this to the special case where $f$ is an affine map with lift $\tilde{f}:G\to F_{n}(G,\pi):x\mapsto(g_{1}\varphi_{1}(x),\ldots,g_{n}\varphi_{n}(x))$. In this case, the formula above can be reduced to an expression purely in terms of the morphisms $\varphi_{i}$.
Recall that, for a morphism $\phi^{\prime}:S^{\prime}\to N$ between lattices in a connected and simply connected nilpotent Lie group $G$, we use $(\phi^{\prime})_{*}:\mathfrak{g}\to\mathfrak{g}$ to denote the Lie algebra morphism induced by the unique extension of $\phi^{\prime}$ to $G\to G$.
Lemma 4.1.
Suppose $\pi\backslash G$ is an infra-nilmanifold finitely covered by a nilmanifold $N\backslash G$, and $S$ is a finite index subgroup of $\pi$. Let $S^{\prime}$ be the subgroup generated by $\{\gamma^{[\pi:N]}\mid\gamma\in S\}$. Let $\phi:S\to\pi$ be a morphism and $\phi^{\prime}:S^{\prime}\to N$ its restriction to $S^{\prime}$. If $\det(I-(\phi^{\prime})_{*})\neq 0$, then $\mathrm{fix}(\phi)=1$.
Proof.
We first prove that $\mathrm{fix}(\phi^{\prime})=1$. To this end, let $\exp:\mathfrak{g}\to G$ denote the exponential map of $G$, and $\log:G\to\mathfrak{g}$ its inverse.
Suppose $\gamma\in S^{\prime}$ is a fixed point of $\phi^{\prime}$. Then $X=\log(\gamma)$ is a fixed point of $(\phi^{\prime})_{*}$, i.e. $(I-(\phi^{\prime})_{*})(X)=0$. Since $\det(I-(\phi^{\prime})_{*})\neq 0$, the only element $X$ satisfying this equation is 00. It follows that $\gamma=\exp(0)=1$.
Now, suppose $\gamma\in S$ is a fixed point of $\phi$, and denote $k=[\pi:N]$. Then $\phi(\gamma^{k})=\phi(\gamma)^{k}=\gamma^{k}$, so $\gamma^{k}\in S^{\prime}$ is a fixed point of $\phi^{\prime}$. By the above, we have $\gamma^{k}=1$. Since $S$ is torsion free, it follows that $\gamma=1$. ∎
We will now show that for an affine map, the number $\varepsilon_{g\alpha,i}$ in (3.1) is independent of $g\in N$, using the results from [4]. We distinguish two cases. In both cases we will need the following lemma.
Lemma 4.2 ([4, Lemma 6.4]).
Let $G$ be a connected and simply connected nilpotent Lie group, and $\varphi\in\mathrm{End}(G)$. If $\varphi_{*}$ has no eigenvalue $1$, the map $G\to G:x\mapsto\varphi(x)x^{-1}$ is bijective.
The first case is the following.
Notation 4.3.
For $\alpha\in\pi\subseteq G\rtimes\mathrm{Aut}(G)$ we will write $\alpha=(a_{\alpha},A_{\alpha})$ with $a_{\alpha}\in G$ and $A_{\alpha}\in\mathrm{Aut}(G)$.
Corollary 4.4.
Let $f:\pi\backslash G\to D_{n}(\pi\backslash G)$ be an affine $n$-valued map with lift $\tilde{f}:x\mapsto(g_{1}\varphi_{1}(x),\ldots,g_{n}\varphi_{n}(x))$ and corresponding numbers $\varepsilon_{g\alpha,i}$. Fix $i$ and $\alpha\in\pi$. If $\det(I-(A_{\alpha}\varphi_{i})_{*})\neq 0$, then $\varepsilon_{g\alpha,i}=1$ for all $g\in N$.
Proof.
Pick $g\in N$. Note that for all $x\in G$,
$g\alpha\tilde{f}_{i}(x)=g\alpha g_{i}\varphi_{i}(x)=ga_{\alpha}A_{\alpha}(g_{i})A_{\alpha}\varphi_{i}(x).$
So if we denote $ga_{\alpha}A_{\alpha}(g_{i})=g^{\prime}$, then $g\alpha\tilde{f}_{i}=g^{\prime}A_{\alpha}\varphi_{i}$.
By Lemma 4.2, the map $G\to G:x\mapsto A_{\alpha}\varphi_{i}(x)x^{-1}$ is a bijection. Hence there is a unique $x\in G$ such that $A_{\alpha}\varphi_{i}(x)x^{-1}=(g^{\prime})^{-1}$, i.e. such that $g^{\prime}A_{\alpha}\varphi_{i}(x)=x$. That is, the map $g\alpha\tilde{f}_{i}=g^{\prime}A_{\alpha}\varphi_{i}$ has a unique fixed point. The index of the fixed point class $p(\mathrm{Fix}(g\alpha\tilde{f}_{i}))$ is then equal to the index of this unique fixed point of $g^{\prime}A_{\alpha}\varphi_{i}$.
In [4, Lemma 6.10] it is proven that, if $\varphi:G\to G$ is a morphism so that $G\to G:x\mapsto\varphi(x)x^{-1}$ is a bijection, then for any $h\in G$, the index of $G\to G:x\mapsto h\varphi(x)$ at its unique fixed point is ±$\pm 1$. Applying this to $\varphi=A_{\alpha}\varphi_{i}$ and $h=g^{\prime}$ gives the desired result. ∎
For the second case, we can use the following lemma, which is a special case of [4, Lemma 6.6]. We give a more polished version of the proof adapted to our specific needs.
Lemma 4.5.
Let $G$ be a connected and simply connected nilpotent Lie group, and $\varphi\in\mathrm{End}(G)$. If $\varphi_{*}$ has eigenvalue $1$, then for all $g\in G$ there is a $g_{0}\in G$ such that $g\varphi(x)g_{0}\neq x$ for all $x\in G$.
Proof.
We prove the claim by induction on the nilpotency class $c$ of $G$.
If $c=1$, then $G=\mathbb{R}^{k}$ for some $k$ and $\varphi=\varphi_{*}$ is a linear map represented by a matrix which has eigenvalue $1$. We can choose a basis $\{v_{1},\ldots,v_{k}\}$ of $\mathbb{R}^{k}$ with respect to which the matrix of $\varphi$ is of the form
$\begin{bmatrix}1&0&\cdots&0\\ *&*&\cdots&*\\ \vdots&\vdots&\ddots&\vdots\\ *&*&\cdots&*\end{bmatrix}.$
That is, if we write $V=\langle v_{2},\ldots,v_{k}\rangle$, then $\varphi(\mu v_{1}+v)\equiv\mu v_{1}\;\mathrm{mod}\,V$ for all $\mu\in\mathbb{R}$ and $v\in V$.
Take $\lambda\in\mathbb{R}$ such that $g\equiv\lambda v_{1}\;\mathrm{mod}\,V$. We claim that $g_{0}=(1-\lambda)v_{1}$ does the job. Indeed, suppose $x\in\mathbb{R}^{k}$ satisfies $g+\varphi(x)+g_{0}=x$. If we write $x=\mu v_{1}+v$ with $v\in V$, this reduces to
$\lambda v_{1}+\mu v_{1}+(1-\lambda)v_{1}=\mu v_{1}\;\mathrm{mod}\,V,$
or equivalentely, $v_{1}\in V$, which is a contradiction.
Now suppose $c>1$ and the result holds for all groups of smaller nilpotency class. For all $i=1,\ldots,c$, let $M_{i}$ be the matrix of the morphism induced by $\varphi$ on $\gamma_{i}(G)/\gamma_{i+1}(G)$. We distinguish two cases: either one of the matrices $M_{1},\ldots,M_{c-1}$ has eigenvalue $1$, or $M_{c}$ is the only matrix with eigenvalue $1$.
In the first case, we can apply the induction hypothesis to the map $G/\gamma_{c}(G)\to G/\gamma_{c}(G):\bar{x}\mapsto\overline{\varphi(x)}$ (and the element $\bar{g}$ for $g$) to obtain a $g_{0}\in G$ such that $\overline{g\varphi(x)g_{0}}\neq\bar{x}$ in $G/\gamma_{c}(G)$, so in particular $g\varphi(x)g_{0}\neq x$ in $G$, for all $x\in G$.
In the second case, we can apply the induction hypothesis to the restriction of $\varphi$ to $\gamma_{c}(G)$ (and the identity element for $g$) to obtain a $g_{0}\in\gamma_{c}(G)$ such that $\varphi(x)g_{0}\neq x$ for all $x\in\gamma_{c}(G)$. On the other hand, we can apply Lemma 4.2 to the map $G/\gamma_{c}(G)\to G/\gamma_{c}(G):\bar{x}\mapsto\overline{\varphi(x)}$, which has no eigenvalue $1$. It follows that there is a unique $\bar{x}_{0}\in G/\gamma_{c}(G)$ such that $\overline{x_{0}\varphi(x_{0})^{-1}}=\bar{g}$. Choose a representative $x_{0}\in G$ for $\bar{x}_{0}$, and take $z\in\gamma_{c}(G)$ such that $g\varphi(x_{0})=x_{0}z$. We claim that the element $g_{0}z^{-1}$ does the job. Indeed, suppose $x\in G$ satisfies $g\varphi(x)g_{0}z^{-1}=x$. By uniqueness of $\bar{x}_{0}$, this element $x$ must be of the form $x_{0}z^{\prime}$ for some $z^{\prime}\in\gamma_{c}(G)$. Then we have
$g\varphi(x_{0}z^{\prime})g_{0}z^{-1}=x_{0}z^{\prime}$ $\enskip\Leftrightarrow\enskip g\varphi(x_{0})\varphi(z^{\prime})g_{0}z^{-1}=x_{0}z^{\prime}$
$\enskip\Leftrightarrow\enskip x_{0}z\varphi(z^{\prime})g_{0}z^{-1}=x_{0}z^{\prime}$
$\enskip\Leftrightarrow\enskip\varphi(z^{\prime})g_{0}=z^{\prime},$
which is impossible for $z^{\prime}\in\gamma_{c}(G)$, by definition of $g_{0}$. ∎
Corollary 4.6.
Let $f:\pi\backslash G\to D_{n}(\pi\backslash G)$ be an affine $n$-valued map with lift $\tilde{f}:x\mapsto(g_{1}\varphi_{1}(x),\ldots,g_{n}\varphi_{n}(x))$ and corresponding numbers $\varepsilon_{g\alpha,i}$. Fix $i$ and $\alpha\in\pi$. If $\det(I-(A_{\alpha}\varphi_{i})_{*})=0$, then $\varepsilon_{g\alpha,i}=0$ for all $g\in N$.
Proof.
Pick $g\in N$ and define $g^{\prime}$ as in Corollary 4.4. By Lemma 4.5, there exits a $g_{0}\in G$ such that $g^{\prime}A_{\alpha}\varphi_{i}(x)g_{0}\neq x$ for all $x\in G$. Then we have a homotopy
$\pi\backslash G\times[0,1]\to D_{n}(\pi\backslash G):(p(x),t)\mapsto\{p(\tilde{f}_{1}(x)g_{0}^{t}),\ldots,p(\tilde{f}_{n}(x)g_{0}^{t})\}.$
Under this homotopy, the fixed point class $p(\mathrm{Fix}(g\alpha\tilde{f}_{i}))$ of $f$ corresponds to the fixed point class $p(\mathrm{Fix}(g\alpha\tilde{f}_{i}g_{0}))$. Since $g\alpha\tilde{f}_{i}g_{0}$ has no fixed points by definition of $g_{0}$, the index of this fixed point class is 00. As indices are preserved under homotopies, the index of $p(\mathrm{Fix}(g\alpha\tilde{f}_{i}))$ is also 00. ∎
On the other hand, the results in [4] give an expression for the number of terms in the second sum in (3.1):
Lemma 4.7.
For a morphism $\phi^{\prime}:S^{\prime}\to N$, with $N$ finitely generated, torsion free and nilpotent, and $S^{\prime}$ a finite index subgroup of $N$, one has
$R(\phi^{\prime})=[N:S^{\prime}]|\det(I-(\phi^{\prime})_{*})|_{\infty},$
where $|a|_{\infty}$ is defined as $|a|$ if $a\neq 0$, and $\infty$ if $a=0$.
Proof.
Let $c$ be the nilpotency class of $N$. For $i=1,\ldots,c$, write $N_{i}=\sqrt[N]{\gamma_{i}(N)}$ and $S^{\prime}_{i}=\sqrt[S^{\prime}]{\gamma_{i}(S^{\prime})}$. If $M_{i}$ denotes the matrix of the morphism $S^{\prime}_{i}N_{i+1}/N_{i+1}\to N_{i}/N_{i+1}$ induced by $\phi^{\prime}$, it is proven in [4, Theorem 4.14] that
$R(\phi^{\prime})=[N:S^{\prime}]|\det(I-M)|_{\infty},$
where
$M=\begin{bmatrix}M_{1}&0&\cdots&0\\ 0&M_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&M_{c}\end{bmatrix}.$
On the other hand, it is proven in [4, Theorem 5.4] that the matrix of $(\phi^{\prime})_{*}$ with respect to a well-chosen basis for $\mathfrak{g}$ is of the form
$\begin{bmatrix}M_{1}&0&\cdots&0\\ *&M_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ *&*&\cdots&M_{c}\end{bmatrix}.$
Thus, we can replace $\det(I-M)$ by $\det(I-(\phi^{\prime})_{*})$, which gives the desired formula. ∎
Lastly, we observe:
Lemma 4.8.
Let $f:\pi\backslash G\to D_{n}(\pi\backslash G)$ be an affine $n$-valued map with lift $\tilde{f}:x\mapsto(g_{1}\varphi_{1}(x),\ldots,g_{n}\varphi_{n}(x))$ and corresponding morphisms $\phi_{i}^{\prime}:S_{i}^{\prime}\to N$. For all $i$ and $\alpha\in\pi$, for all $g\in N$,
$\det(I-(\tau_{g\alpha}\phi^{\prime}_{i})_{*})=\det(I-(A_{\alpha}\varphi_{i})_{*}).$
Proof.
We may choose a basis $v_{1,1},v_{1,2},\ldots,v_{1,k_{1}},v_{2,1},\ldots,v_{c,k_{c}}$ for $\mathfrak{g}$ in such a way that the vectors $v_{i,1},v_{i,2},\ldots,v_{i,k_{i}},v_{i+1,1},\ldots,v_{c,k_{c}}$ form a basis of $\gamma_{i}(\mathfrak{g})$, the $i$-th term in the lower central series of the Lie algebra $\mathfrak{g}$. With respect to this basis, all morphisms $\mathfrak{g}\to\mathfrak{g}$ are represented by blocked lower triangular matrices (with blocks of size $k_{i}\times k_{i}$ on the diagonal).
Now note that for the morphisms $\tau_{g\alpha},\,\phi_{i}^{\prime}:S_{i}^{\prime}\to N$, and therefore also for their extensions $G\to G$,
$\tau_{g\alpha}=c_{ga_{\alpha}}A_{\alpha}\qquad\text{and}\qquad\phi_{i}^{\prime}=c_{g_{i}}\varphi_{i}$
where $c_{g}:G\to G$ denotes conjugation with $g\in G$. (The equality for $\phi_{i}$ follows by evaluating (2.1) at $x=1$ and $\gamma\in S_{i}^{\prime}$.)
For all $g\in G$, the map $c_{g}:G\to G$ induces the identity on $\gamma_{i}(G)/\gamma_{i+1}(G)$, so the maps $(c_{g})_{\ast}$ induce the identity on $\gamma_{i}(\mathfrak{g})/\gamma_{i+1}(\mathfrak{g})$ and the matrix representation of $(c_{g})_{\ast}$ has identity blocks on the diagonal. It follows that the matrices of $(\tau_{g\alpha})_{*}$ and $(\phi^{\prime}_{i})_{*}$ have the same blocks on the diagonal as those of $(A_{\alpha})_{*}$ and $(\varphi_{i})_{*}$, respectively; therefore
$\det(I-(\tau_{g\alpha}\phi^{\prime}_{i})_{*})=\det(I-(A_{\alpha}\varphi_{i})_{*}).\qed$
Putting all these results together gives the averaging formula:
Theorem 4.9.
The Nielsen number of an affine map $f:\pi\backslash G\to D_{n}(\pi\backslash G)$ with lift $\tilde{f}:x\mapsto(g_{1}\varphi_{1}(x),\ldots,g_{n}\varphi_{n}(x))$ is given by
$N(f)=\frac{1}{[\pi:N]}\sum_{\bar{\alpha}\in\pi/N}\sum_{i=1}^{n}|\det(I-(A_{\alpha}\varphi_{i})_{*})|.$
Proof.
Recall from section 3 that
$N(f)=\sum_{i=1}^{n}\frac{1}{[\pi:S^{\prime}_{i}]}\sum_{\bar{\alpha}\in\pi/N}\sum_{[g]\in\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]}|\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}|\,\varepsilon_{g\alpha,i}.$
In case $\det(I-(A_{\alpha}\varphi_{i})_{*})\neq 0$, Lemma 4.1 implies $\mathrm{fix}(\tau_{g\alpha}\phi_{i})=1$, so in particular $|\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}|=1$; Corollary 4.4 implies $\varepsilon_{g\alpha,i}=1$ for all $g$; and Lemmas 4.7 and 4.8 imply
$R(\tau_{\alpha}\phi^{\prime}_{i})=[N:S^{\prime}_{i}]|\det(I-(A_{\alpha}\varphi_{i})_{*})|.$
Thus, in this case,
$\sum_{[g]\in\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]}|\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}|\,\varepsilon_{g\alpha,i}=[N:S^{\prime}_{i}]|\det(I-(A_{\alpha}\varphi_{i})_{*})|.$
In case $\det(I-(A_{\alpha}\varphi_{i})_{*})=0$, Corollary 4.6 implies $\varepsilon_{g\alpha,i}=0$ for all $g$, so
$\sum_{[g]\in\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]}|\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}|\,\varepsilon_{g\alpha,i}=0=[N:S^{\prime}_{i}]|\det(I-(A_{\alpha}\varphi_{i})_{*})|.$
It follows that
$N(f)$ $=\sum_{i=1}^{n}\frac{1}{[\pi:S^{\prime}_{i}]}\sum_{\bar{\alpha}\in\pi/N}[N:S^{\prime}_{i}]|\det(I-(A_{\alpha}\varphi_{i})_{*})|$
$=\frac{1}{[\pi:N]}\sum_{\bar{\alpha}\in\pi/N}\sum_{i=1}^{n}\,|\det(I-(A_{\alpha}\varphi_{i})_{*})|.\qed$
5. Example
To illustrate the formula, consider the following example. Take $G=\mathbb{R}^{2}$ and let $\pi\subseteq\mathbb{R}^{2}\rtimes\mathrm{GL}_{2}(\mathbb{R})$ be the group generated by
$a=\left(\begin{bmatrix}1\\ 0\end{bmatrix},\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\right),\quad b=\left(\begin{bmatrix}0\\ \frac{1}{2}\end{bmatrix},\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\right).$
Then $\pi\backslash\mathbb{R}^{2}$ is an infra-nilmanifold isomorphic to the Klein bottle. Consider the map given by
$\tilde{f}:\mathbb{R}^{2}\to F_{2}(\mathbb{R}^{2},\pi):(t_{1},t_{2})\mapsto((0,\textstyle\frac{t_{2}}{2}),(0,\textstyle\frac{t_{2}}{2}-\frac{1}{4})).$
This map induces a well-defined $2$-valued map $f:\pi\backslash\mathbb{R}^{2}\to D_{2}(\pi\backslash\mathbb{R}^{2})$ on the Klein bottle, since for all $(t_{1},t_{2})\in\mathbb{R}^{2}$
$\tilde{f}(a(t_{1},t_{2}))$ $=((0,\textstyle\frac{t_{2}}{2}),(0,\textstyle\frac{t_{2}}{2}-\frac{1}{4}))=(1,1;1)\tilde{f}(t_{1},t_{2})$
$\tilde{f}(b(t_{1},t_{2}))$ $=(b(0,\textstyle\frac{t_{2}}{2}-\frac{1}{4}),(0,\textstyle\frac{t_{2}}{2}))=(b,1;\sigma)\tilde{f}(t_{1},t_{2})$
with $\sigma\in\Sigma_{2}$ the permutation that interchanges $1$ and $2$.
Note that $f$ is an affine map, with
$g_{1}=\begin{bmatrix}0\\ 0\end{bmatrix},\quad g_{2}=\begin{bmatrix}0\\ -\frac{1}{4}\end{bmatrix},\quad\varphi_{1}=\varphi_{2}=\begin{bmatrix}0&0\\ 0&\frac{1}{2}\end{bmatrix}.$
To compute the Nielsen number of $f$, we consider the nilmanifold $Z\backslash\mathbb{R}^{2}$ where $Z=\pi\cap\mathbb{R}^{2}=\langle a,b^{2}\rangle\cong\mathbb{Z}^{2}$, i.e. the $2$-torus. The group $\pi/Z$ has two elements, represented by
$\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\right)\quad\text{and}\quad\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\right).$
Thus, by Theorem 4.9
$N(f)$ $=\frac{1}{2}\sum_{\bar{\alpha}\in\pi/Z}\sum_{i=1}^{2}\left|\,\det\left(I-A_{\alpha}\begin{bmatrix}0&0\\ 0&\frac{1}{2}\end{bmatrix}\right)\right|$
$=\sum_{\bar{\alpha}\in\pi/Z}\left|\,\det\left(I-A_{\alpha}\begin{bmatrix}0&0\\ 0&\frac{1}{2}\end{bmatrix}\right)\right|$
$=\left|\,\det\left(I-\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\begin{bmatrix}0&0\\ 0&\frac{1}{2}\end{bmatrix}\right)\right|+\left|\,\det\left(I-\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\begin{bmatrix}0&0\\ 0&\frac{1}{2}\end{bmatrix}\right)\right|$
$=\left|\,\det\begin{bmatrix}1&0\\ 0&\frac{1}{2}\end{bmatrix}\right|+\left|\,\det\begin{bmatrix}1&0\\ 0&\frac{1}{2}\end{bmatrix}\right|=1.$
A direct calculation shows that the map $f$ has one isolated fixed point: indeed, for all $k,\ell\in\mathbb{Z}$ we have
$(t_{1},t_{2})=a^{k}b^{\ell}(0,\textstyle\frac{t_{2}}{2})$ $\enskip\Leftrightarrow\enskip(t_{1},t_{2})=(k,\textstyle\frac{t_{2}}{2}+\frac{\ell}{2})$
$\enskip\Leftrightarrow\enskip(t_{1},t_{2})=(k,\ell)$
and
$(t_{1},t_{2})=a^{k}b^{\ell}(0,\textstyle\frac{t_{2}}{2}-\frac{1}{4})$ $\enskip\Leftrightarrow\enskip(t_{1},t_{2})=(k,\textstyle\frac{t_{2}}{2}-\frac{1}{4}+\frac{\ell}{2})$
$\enskip\Leftrightarrow\enskip(t_{1},t_{2})=(k,\textstyle-\frac{1}{2}+\ell);$
and all these points $(k,\ell)$ and $(k,-\frac{1}{2}+\ell)$ in $\mathbb{R}^{2}$ project to the same point on the Klein bottle. Thus, the minimal number of fixed points is actually attained in this case.
Remark 5.1.
This map $f$ is an example of a map on the Klein bottle that does not lift to the torus $Z\backslash\mathbb{R}^{2}$. In fact, there is no nilmanifold (torus) $N\backslash\mathbb{R}^{2}$ that finitely covers $\pi\backslash\mathbb{R}^{2}$ so that $f$ lifts to $N\backslash\mathbb{R}^{2}$: in order for such a lift to exist, the induced morphism of covering groups $f_{*}=(\phi_{1},\phi_{2};\sigma):\pi\to\pi^{2}\rtimes\Sigma_{2}$ defined in section 1 must satisfy $f_{*}(N)\subseteq N^{2}\rtimes\Sigma_{2}$. However, note that $f_{*}(b^{2})=(b,b;1)$. If $N\backslash\mathbb{R}^{2}$ is a nilmanifold that finitely covers $\pi\backslash\mathbb{R}^{2}$, then $N$ must be a finite index subgroup of $Z$. Hence there is a smallest $\ell\in\mathbb{Z}_{>0}$ so that $(b^{2})^{\ell}\in N$. But then $f_{*}((b^{2})^{\ell})=(b^{\ell},b^{\ell};1)$ is not an element of $N^{2}\rtimes\Sigma_{2}$, by minimality of $\ell$.
This illustrates the fact that a classical averaging formula in terms of lifts as in [8] is not feasible in the $n$-valued case.